a)Price of Pepsi is Rs.120 and its quantity demanded is 300 bottles. Suppose due to Raman Pepsi company reduces price of it to Rs. 108 ,due to it its demand increases to 360 bottles. Calculate elasticity and also mention the type of elasticity.
1
Expert's answer
2020-04-29T09:21:23-0400

Elasticity of demand refers to the responsive change of quantity demanded as a result of changes in factors affecting the demand. One of the factor that influence the demand is the price.

The Price elasticity of demand is calculated using the following formula:

"\text{Arc price elasticity of demand}=\dfrac{\Delta Q}{\Delta P}*\dfrac{\bar{P}}{\bar {Q} }"

"\bar{P} \bar {Q}" are the price and the Quantity mid point.

"\text{Arc price elasticity of demand}=\dfrac{360-300}{108-120}*\dfrac{ \dfrac{120+108}{2}}{\dfrac{360+300}{2}}"


"\text{Arc price elasticity of demand}=\dfrac{60}{-12}* \dfrac{114}{330}"


"\text{Arc price elasticity of demand}=-1.73"


Therefore, since the elasticity is more than 1 in absolute terms, then we can conclude that the type of elasticity is elastic in nature.